Using the following chemical equation how many grams of carbon dioxide can be produced by burning 7.99 g of oxygen?

C4H10 + 5O2 -> 3CO2 + 4H2O


1
Expert's answer
2021-02-15T02:47:23-0500

C4H10 + 5O2 -> 3CO2 + 4H2O

According to the equation, the mass of carbon dioxide equals:

m (CO2) = m(O2) × 3 × Mr(CO2) / [5 × Mr(O2)] = 7.99 g × 3 × 44 g/mol / [5 × 32 g/mol] = 6.6 g


Answer: 6.6 g